solve.
Juan goes to a bank and gets change for a $50 bill consisting of $5 bills and $1 bills. There are 22 bills in all. How many of each kind are there?

Found 2 solutions by checkley71, stanbon:
Answer by checkley71(8403)

F+O=22 OR F=22-0
5F+10=50
5(22-O)+10=50
110-5O+O=50
-4O=50-110
-4O=-60
O=-60/-4
O=15 NUMBER OF ONE DOLLAR BILLS
22-15=7 FIVE DOLLAR BILLS.
PROOF:
5*7+15*1=50
35+15=50
50=50

Answer by stanbon(48575)

Juan goes to a bank and gets change for a $50 bill consisting of $5 bills and $1 bills. There are 22 bills in all. How many of each kind are there?
---------------
Let the number of $1 bills be "x" ; value is x dollars
-------------------
# of $5 bills is "22-x" ; value is 5*(22-x)= 110-5x dollars
---------------------
EQUATION:
value + value = 50 dollars
x + 110-5x = 50
-4x = -60
x = 15 (# of $1 bills)
22-x = 7 (# of $5 bills)
============================
Cheers
Stan H.